# Manifold-Constrained Nucleus-Level Denoising Diffusion Model for Structure-Based Drug Design

### v1

@@ 1 Introduction
Our Contributions. To tackle this challenge, we propose NucleusDiff, a manifold-constrained denoising diffusion model for structure-based drug design. To incorporate inductive biases from quantum physics, NucleusDiff directly models the constraints between atomic nuclei and their surrounding electron clouds to prevent separation violation. In essence, NucleusDiff employs two sub-modules: one for modeling the atomic nuclei and the other one for modeling the manifold, a sphere corresponding to the average distance (a.k.a. atomic radii) from the nuclei to the outermost electrons in the electron cloud. Then NucleusDiff leverages a regularization term to align the distance between nuclei and sampled manifold points with the atomic radii. This alignment implicitly maintains proper pairwise atomic distances and upholds the principles of attraction and repulsion that govern molecular interaction. Additionally, the complexity of modeling such a manifold constraint increases linearly to the number of atoms, making NucleusDiff an efficient paradigm.
Figure 1: (a) Illustration of the nucleus, the electron cloud, and the manifold of an atom. The electron cloud represents the probabilistic distribution of electrons around the nucleus, and the manifold is the sphere corresponding to the average distance from the nucleus to the outermost electrons in the electron cloud. (b) Illustration of the manifold surrounding a molecule. (c) Illustration of the mesh points obtained from discretizing a manifold. (d) Pipeline of NucleusDiff. NucleusDiff performs denoising diffusion on both the nuclei and the discretized mesh points, where the distances between them approximate the atomic radii.
@@ 2.1 Background
Each atom constitutes a nucleus and an electron cloud surrounds each nucleus, as shown in Figure 1. Recent works employ manifold learning on such an electron cloud for molecule property prediction [10, 11] and protein modeling in structure-based drug design [12]. In this work, we consider modeling the manifold around each nucleus in the ligands, and the approximated distance is the atomic radii. Then we discretize the manifold into triangle mesh points, a form suitable for computational analysis. This is implemented using the Python package PyMesh [13]. For notation, for each ligand atom $({\bm{v}}^{L},{\bm{x}}^{L})$, the coordinate and type of nuclei are the same as the atom-level, i.e., ${\bm{v}}^{N}\triangleq{\bm{v}}^{L}$ and ${\bm{x}}^{N}\triangleq{\bm{x}}^{L}$. The coordinate of the discretized points on the manifold is marked as ${\bm{x}}^{M}$. Notice that we use a special token to delegate the electron points on the manifold.
@@ 2.2 Separation Violation and Violation Ratio Metrics
The separation violation occurs when two atoms come too close to each other, contradicting the underlying physical laws. We introduce the van der Waals radii $d$ for measuring it. Suppose we have one ligand atom coordinate ${\bm{x}}_{i}$, one protein atom coordinate ${\bm{x}}_{j}$, and the corresponding van der Waals radii are $d_{i}$ and $d_{j}$, respectively. Then during the sampling stage for ligand generation, if two atoms get too close to each other, i.e., $\|{\bm{x}}_{j}-{\bm{x}}_{j}\|\leq D_{ij}=d_{i}+d_{j}$, we refer to this as the separation violation problem. To quantify this in existing deep generative models, we propose three violation ratio metrics from three granularities. For clarity, we only present one of the representative metrics in the main article.
@@ 2.3 Manifold-Constrained Nucleus-Level DDPM: NucleusDiff
To reduce the separation violation, we introduce an extra constraint term that keeps the distance between the nucleus and the manifold over the electron cloud as atomic radii, $R$. This constraint aligns with the fundamental physical principles, as electron clouds exert both attraction and repulsion forces, preventing separation violation. To adopt this into modeling, for each nucleus, we obtain its $K$ closest mesh points in the manifold, marked as ${\bm{x}}^{M}$. Thus, the goal becomes to maximize the joint distribution of nuclei and mesh points conditioned on the pocket, as $p({\bm{v}}^{N},{\bm{x}}^{N},{\bm{x}}^{M}|{\bm{v}}^{P},{\bm{x}}^{P})$. To adapt this into Equation 3, the objective of the sampled mesh points is
@@
On the other hand, recall that the mesh points are scattered around the nuclei with a fixed atomic radius $R$. Thus we add a regularization term by forcing the distance between each corresponding mesh point ${\bm{x}}^{N}_{j}$ and nucleus ${\bm{x}}^{M}_{i}$ to be close to $R$:
@@ 2.4 Experimental Setup
We utilize MSMS [17] to compute the solvent-excluded surface of the molecule. It generates a triangular mesh data structure for small molecules, and we set the argument probe radius to 1.5 Å and sampling density argument to 3.0. MSMS possesses certain challenges such as degenerate vertices and disconnected surfaces, which can disrupt the uniform distribution of mesh points. We employ PyMesh to overcome these issues [13]. It enhances the mesh quality by reducing the vertex count and correcting errors in poorly meshed regions. Finally, we select the $K$ mesh points that are closest to the Van der Waals radii distance from the nucleus to construct a mesh point dataset. Notably, this mesh point data structure, which predominantly includes the 3D coordinates of the mesh points, is only required during the learning phase and not during inference.
@@ 2.8 The Results for Minimum Distance Constraint
Minimum Distance Constraint (Parallelogram): For protein-ligand atom pairs $(a,b)$ with separation violation, we first identify atom $c$ within the ligand that is closest to ligand atom $b$. Clearly, the line connecting the protein-ligand pair in 3D space intersects with a sphere centered at ligand atom $c$, with a radius equal to the distance between $c$ and $b$. One intersection point obviously becomes $b$, while the other intersection point $b^{\prime}$ becomes the corrected position of $b$ after applying the minimum distance constraint.
Minimum Distance Constraint (Circle): For a protein-ligand pair $(a,b)$ exhibiting separation violation, we first identify the atom $c$ within the ligand that is closest to ligand atom $b$. We ensure that the distance between $a$ and $b$ is less than that between $a$ and $c$. Based on this condition, we construct two spheres: one centered at protein atom $a$ with a radius equal to the van der Waals radius of $a$ and $b$ (noting that the distance between $(a,b)$ is less than the summation of their van der Waals radius), and another centered at ligand atom $c$ with a radius equal to the distance between $c$ and $b$. If these two spheres intersect, their intersection forms a circular region. By deriving the analytical expression for this circle, we sample a new point $b^{\prime}$ on this circle using a predetermined random seed (42). This $b^{\prime}$ effectively avoids separation violations while preserving the ligand’s geometric characteristics. In the case where the two spheres are tangent, the point of tangency serves as the unique corrected position $b^{\prime}$, avoiding separation violations and maintaining the ligand’s geometric properties.
@@ 4.2 Manifold-Constraint Denoising Diffusion Model
To be more concrete, for each nucleus, we construct a discrete manifold, where the radius is the atomic radii $R$. Then for each nucleus, we obtain its $c$ closest mesh points in the manifold, marked as ${\bm{x}}^{M}$. Thus, instead of $p({\bm{v}}^{N},{\bm{x}}^{N}|{\bm{v}}^{P},{\bm{x}}^{P})$, the objective is to maximize the following likelihood $p({\bm{v}}^{N},{\bm{x}}^{N},{\bm{x}}^{M}|{\bm{v}}^{P},{\bm{x}}^{P})$. The objective on the manifold at time $t$ is Equation 5.
On the other hand, recall that the mesh points are scattered around the nuclei with atomic radii $R$. Motivated by this, we add a regularization term by forcing the distance between each mesh point and nuclei to be close to $R$ as in Equation 6.
@@ Appendix B Formulation and Analysis of Proposed Separation Violation Metrics
The phenomenon of separation violation arises when two atoms come sufficiently close to each other such that their electron clouds overlap. This overlap violates the Pauli exclusion principle and the electrostatic repulsion between electrons. To quantify this proximity, we employ van der Waals radii, denoted as $d$. Consider a scenario involving a ligand atom with coordinates ${\bm{x}}_{i}$ and a protein atom with coordinates ${\bm{x}}_{j}$. The respective van der Waals radii for these atoms are $d_{i}$ and $d_{j}$. During the ligand generation sampling process, if the distance between these two atoms becomes excessively small, specifically if $\|{\bm{x}}_{i}-{\bm{x}}_{j}\|\leq D_{ij}=d_{i}+d_{j}$, we classify this event as an separation violation. To systematically analyze and understand this phenomenon, we propose three distinct metrics for its quantification, each derived from different levels of assessment granularity. These metrics provide a comprehensive framework for evaluating the occurrence and implications of separation violations in molecular simulations and structural biology studies.
@@
  • $D_{ij}$: The sum of the van der Waals radii of the $i$-th ligand atom and the $j$-th protein atom.
@@
The second metric we propose is the Atom-Level Violation Ratio (ALVR). This metric provides a more aggregated view by focusing on individual ligand atoms and determining whether they are involved in any violations with protein atoms. The ALVR measures the proportion of ligand atoms that have at least one of their $K$ nearest protein atoms within a critical distance. This critical distance is defined as less than the sum of their respective van der Waals radii.
@@
  • $D_{ij}$: The sum of the van der Waals radii of the $i$-th ligand atom and the $j$-th protein atom.
@@
  • $D_{ij}$: The sum of the van der Waals radii of the $i$-th ligand atom and the $j$-th protein atom.
@@ G.4 The Violation Results for Minimum Distance Constraint
Minimum Distance Constraint (Circle): For a protein-ligand atom pair $(a,b)$ that violates separation constraints, we begin by identifying the ligand atom $c$ closest to atom $b$. We ensure that the distance between atoms $a$ and $b$ is less than that between $a$ and $c$. Based on this condition, we construct two spheres: the first is centered at protein atom $a$ with a radius equal to the sum of the van der Waals radii of atoms $a$ and $b$, recognizing that the distance between $(a,b)$ is less than their combined van der Waals radii. The second sphere is centered at ligand atom $c$ with a radius equal to the distance between atoms $c$ and $b$. If these two spheres intersect, their intersection forms a circular region. Using an analytical expression for this circle, we sample a new position $b^{\prime}$ on the circle, guided by a predetermined random seed (42). This corrected position $b^{\prime}$ resolves the separation violation while preserving the ligand’s geometric characteristics. In cases where the two spheres are tangent, the point of tangency serves as the unique corrected position of $b^{\prime}$, thereby avoiding separation violations and maintaining the structural integrity of the ligand.

### v2

@@ 1 Introduction
Our Contributions. To tackle this challenge, we propose NucleusDiff, a manifold-constrained denoising diffusion model for structure-based drug design. To incorporate inductive biases from quantum physics, NucleusDiff directly models the constraints between atomic nuclei and their surrounding electron clouds to prevent separation violation. In essence, NucleusDiff employs two sub-modules: one for modeling the atomic nuclei and the other one for modeling the manifold, a sphere corresponding to the average distance (a.k.a. van der Waals radius) from the nuclei to the outermost electrons in the electron cloud. Then NucleusDiff leverages a regularization term to align the distance between nuclei and sampled manifold points with the van der Waals radii. This alignment implicitly maintains proper pairwise atomic distances and upholds the principles of attraction and repulsion that govern molecular interaction. Additionally, the complexity of modeling such a manifold constraint increases linearly to the number of atoms, making NucleusDiff an efficient paradigm.
Figure 1: (a) Illustration of the nucleus, the electron cloud, and the manifold of an atom. The electron cloud represents the probabilistic distribution of electrons around the nucleus, and the manifold is the sphere corresponding to the average distance from the nucleus to the outermost electrons in the electron cloud. (b) Illustration of the manifold surrounding a molecule. (c) Illustration of the mesh points obtained from discretizing a manifold. (d) Pipeline of NucleusDiff. NucleusDiff performs denoising diffusion on both the nuclei and the discretized mesh points, where the distances between them approximate the van der Waals radii.
@@ 2.1 Background
Each atom constitutes a nucleus and an electron cloud surrounds each nucleus, as shown in Figure 1. Recent works employ manifold learning on such an electron cloud for molecule property prediction [10, 11] and protein modeling in structure-based drug design [12]. In this work, we consider modeling the manifold around each nucleus in the ligands, and the approximated distance is the van der Waals radii. Then we discretize the manifold into triangle mesh points, a form suitable for computational analysis. This is implemented using the Python package PyMesh [13]. For notation, for each ligand atom $({\bm{v}}^{L},{\bm{x}}^{L})$, the coordinate and type of nuclei are the same as the atom-level, i.e., ${\bm{v}}^{N}\triangleq{\bm{v}}^{L}$ and ${\bm{x}}^{N}\triangleq{\bm{x}}^{L}$. The coordinate of the discretized points on the manifold is marked as ${\bm{x}}^{M}$. Notice that we use a special token to delegate the electron points on the manifold.
@@ 2.2 Separation Violation and Violation Ratio Metrics
The separation violation occurs when two atoms come too close to each other, contradicting the underlying physical laws. We introduce the covalent radius $d$ to measure it, as the covalent radius is a more strict quantity involving the covalent bonding. Suppose we have one ligand atom coordinate ${\bm{x}}_{i}$, one protein atom coordinate ${\bm{x}}_{j}$, and the corresponding covalent radii are $d_{i}$ and $d_{j}$, respectively. Then during the sampling stage for ligand generation, if two atoms get too close to each other, i.e., $\|{\bm{x}}_{j}-{\bm{x}}_{j}\|\leq D_{ij}=d_{i}+d_{j}$, we refer to this as the separation violation problem. To quantify this in existing deep generative models, we propose three violation ratio metrics from three granularities. For clarity, we only present one of the representative metrics in the main article.
@@ 2.3 Manifold-Constrained Nucleus-Level DDPM: NucleusDiff
To reduce the separation violation, we introduce an extra constraint term that keeps the distance between the nucleus and the manifold over the electron cloud as van der Waals radius, $R$. This constraint aligns with the fundamental physical principles, as electron clouds exert both attraction and repulsion forces, preventing separation violation. To adopt this into modeling, for each nucleus, we obtain its $K$ closest mesh points in the manifold, marked as ${\bm{x}}^{M}$. Thus, the goal becomes to maximize the joint distribution of nuclei and mesh points conditioned on the pocket, as $p({\bm{v}}^{N},{\bm{x}}^{N},{\bm{x}}^{M}|{\bm{v}}^{P},{\bm{x}}^{P})$. To adapt this into Equation 3, the objective of the sampled mesh points is
@@
On the other hand, recall that the mesh points are scattered around the nuclei with a fixed van der Waals radius $R$. Thus we add a regularization term by forcing the distance between each corresponding mesh point ${\bm{x}}^{N}_{j}$ and nucleus ${\bm{x}}^{M}_{i}$ to be close to $R$:
@@ 2.4 Experimental Setup
We utilize MSMS [17] to compute the solvent-excluded surface of the molecule. It generates a triangular mesh data structure for small molecules, and we set the argument probe radius to 1.5 Å and sampling density argument to 3.0. MSMS possesses certain challenges such as degenerate vertices and disconnected surfaces, which can disrupt the uniform distribution of mesh points. We employ PyMesh to overcome these issues [13]. It enhances the mesh quality by reducing the vertex count and correcting errors in poorly meshed regions. Finally, we select the $K$ mesh points that are closest to the van der Waals radius from the nucleus to construct a mesh point dataset. Notably, this mesh point data structure, which predominantly includes the 3D coordinates of the mesh points, is only required during the learning phase and not during inference.
@@ 2.8 The Results for Minimum Distance Constraint
Minimum Distance Constraint (Parallelogram): For protein-ligand atom pairs $(a,b)$ with separation violation, we first identify atom $c$ within the ligand that is closest to ligand atom $b$. Clearly, the line connecting the protein-ligand pair in 3D space intersects with a sphere centered at ligand atom $c$, with the covalent radius equal to the distance between $c$ and $b$. One intersection point obviously becomes $b$, while the other intersection point $b^{\prime}$ becomes the corrected position of $b$ after applying the minimum distance constraint.
Minimum Distance Constraint (Circle): For a protein-ligand pair $(a,b)$ exhibiting separation violation, we first identify the atom $c$ within the ligand that is closest to ligand atom $b$. We ensure that the distance between $a$ and $b$ is less than that between $a$ and $c$. Based on this condition, we construct two spheres: one centered at protein atom $a$ with a radius equal to the covalent radius of $a$ and $b$ (noting that the distance between $(a,b)$ is less than the summation of their covalent radius), and another centered at ligand atom $c$ with a radius equal to the distance between $c$ and $b$. If these two spheres intersect, their intersection forms a circular region. By deriving the analytical expression for this circle, we sample a new point $b^{\prime}$ on this circle using a predetermined random seed (42). This $b^{\prime}$ effectively avoids separation violations while preserving the ligand’s geometric characteristics. In the case where the two spheres are tangent, the point of tangency serves as the unique corrected position $b^{\prime}$, avoiding separation violations and maintaining the ligand’s geometric properties.
@@ 4.2 Manifold-Constraint Denoising Diffusion Model
To be more concrete, for each nucleus, we construct a discrete manifold, where the radius is the van der Waals radius $R$. Then for each nucleus, we obtain its $c$ closest mesh points in the manifold, marked as ${\bm{x}}^{M}$. Thus, instead of $p({\bm{v}}^{N},{\bm{x}}^{N}|{\bm{v}}^{P},{\bm{x}}^{P})$, the objective is to maximize the following likelihood $p({\bm{v}}^{N},{\bm{x}}^{N},{\bm{x}}^{M}|{\bm{v}}^{P},{\bm{x}}^{P})$. The objective on the manifold at time $t$ is Equation 5.
On the other hand, recall that the mesh points are scattered around the nuclei with van der Waals radius $R$. Motivated by this, we add a regularization term by forcing the distance between each mesh point and nuclei to be close to $R$ as in Equation 6.
@@ Appendix B Formulation and Analysis of Proposed Separation Violation Metrics
The phenomenon of separation violation arises when two atoms come sufficiently close to each other such that their electron clouds overlap. This overlap violates the Pauli exclusion principle and the electrostatic repulsion between electrons. To quantify this proximity, we employ covalent radius, denoted as $d$. Consider a scenario involving a ligand atom with coordinates ${\bm{x}}_{i}$ and a protein atom with coordinates ${\bm{x}}_{j}$. The respective covalent radii for these atoms are $d_{i}$ and $d_{j}$. During the ligand generation sampling process, if the distance between these two atoms becomes excessively small, specifically if $\|{\bm{x}}_{i}-{\bm{x}}_{j}\|\leq D_{ij}=d_{i}+d_{j}$, we classify this event as an separation violation. To systematically analyze and understand this phenomenon, we propose three distinct metrics for its quantification, each derived from different levels of assessment granularity. These metrics provide a comprehensive framework for evaluating the occurrence and implications of separation violations in molecular simulations and structural biology studies.
@@
  • $D_{ij}$: The sum of the covalent radius of the $i$-th ligand atom and the $j$-th protein atom.
@@
The second metric we propose is the Atom-Level Violation Ratio (ALVR). This metric provides a more aggregated view by focusing on individual ligand atoms and determining whether they are involved in any violations with protein atoms. The ALVR measures the proportion of ligand atoms that have at least one of their $K$ nearest protein atoms within a critical distance. This critical distance is defined as less than the sum of their respective covalent radii.
@@
  • $D_{ij}$: The sum of the covalent radius of the $i$-th ligand atom and the $j$-th protein atom.
@@
  • $D_{ij}$: The sum of the covalent radius of the $i$-th ligand atom and the $j$-th protein atom.
@@ G.4 The Violation Results for Minimum Distance Constraint
Minimum Distance Constraint (Circle): For a protein-ligand atom pair $(a,b)$ that violates separation constraints, we begin by identifying the ligand atom $c$ closest to atom $b$. We ensure that the distance between atoms $a$ and $b$ is less than that between $a$ and $c$. Based on this condition, we construct two spheres: the first is centered at protein atom $a$ with a radius equal to the sum of the covalent radius of atoms $a$ and $b$, recognizing that the distance between $(a,b)$ is less than their combined covalent radius. The second sphere is centered at ligand atom $c$ with a radius equal to the distance between atoms $c$ and $b$. If these two spheres intersect, their intersection forms a circular region. Using an analytical expression for this circle, we sample a new position $b^{\prime}$ on the circle, guided by a predetermined random seed (42). This corrected position $b^{\prime}$ resolves the separation violation while preserving the ligand’s geometric characteristics. In cases where the two spheres are tangent, the point of tangency serves as the unique corrected position of $b^{\prime}$, thereby avoiding separation violations and maintaining the structural integrity of the ligand.